# Threshold Filtering Packing for Supervised Fine-Tuning: Training Related Samples within Packs

### v1

@@ Abstract
Packing for Supervised Fine-Tuning (SFT) in autoregressive models involves concatenating data points of varying lengths until reaching the designed maximum length to facilitate GPU processing. However, randomly concatenating data points and feeding them into an autoregressive transformer can lead to cross-contamination of sequences due to the significant difference in their subject matter. The mainstream approaches in SFT ensure that each token in the attention calculation phase only focuses on tokens within its own short sequence, without providing additional learning signals for the preceding context. To address these challenges, we introduce Threshold Filtering Packing (TFP), a method that selects samples with related context while maintaining sufficient diversity within the same pack. Our experiments show that TFP offers a simple-to-implement and scalable approach that significantly enhances SFT performance, with observed improvements of up to 7% on GSM8K, 4% on HumanEval, and 15% on the adult-census-income dataset.
@@
Jiancheng Dong1, Lei Jiang1, Wei Jin2, Lu Cheng1 1University of Illinois Chicago, 2Emory University dongjiancheng77@gmail.com, ljian43@uic.edu, wei.jin@emory.edu, lucheng@uic.edu
@@ 1 Introduction
In Supervised Fine-Tuning (SFT) for large language models (LLMs), sequence lengths can vary substantially, requiring the wrapping of data into tensors using PyTorch to apply matrix operations optimized for CUDA and GPUs Raffel et al. (2023). As illustrated in Figure 1(a), vanilla fine-tuning pad shorter sequences with special tokens "[PAD]" up to the maximum sequence length. While this ensures uniformity, it introduces inefficiencies by including irrelevant padding tokens in the computation, wasting GPU resources and dilutes the model’s learning signal Kundu et al. (2024).
@@
However, randomly concatenating these data samples, which are then fed into an autoregressive transformer (Figure 1(b)), can result in sequence cross-contamination Krell et al. (2022). This occurs when predictions for one sequence are influenced by an unrelated sequence, making accurate predictions challenging, especially when the sequences pertain to different subjects. An example of cross-contamination is when a model is instructed to generate a multiplication table, followed by an instruction to create a list of useful expressions in French. If these two sequences are concatenated without careful separation, the model might incorrectly mix instructions and outputs, such as generating multiplication results interspersed with French phrases like "3 x 2 = Bonjour," leading to inappropriate and confusing outcomes. Moreover, current SFT pipelines Kundu et al. (2024) cause previous samples to provide no signal for predicting the next sample, thereby reducing learning efficiency and negatively impacting the few-shot performance of LLMs.
To address these challenges, in this work, we present Threshold Filtering Packing (TFP), a new packing approach that packs sequences of related yet diverse samples, encouraging context richness and reasoning across sample boundaries. Specifically, we employ a greedy algorithm inspired by the Traveling Salesman Problem (TSP) Applegate et al. (2006) to efficiently map out a path for segmentation into multiple packs. TFP is implemented to further refine these packs by ensuring that overly similar samples are not grouped together. Setting the threshold is crucial in this context as it allows us to strike a balance between similarity and diversity within each pack, preventing homogeneity and ensuring the robustness of the generated packs. As shown in Figure 1(c), each sample is first converted to an embedding and then represented as a node in the graph. As shown in Figure 1(d), TFP employs threshold filtering to ensure each sample is distinct enough from recent ones, preventing the model from merely replicating previous outputs. This method results in packs that provide useful context while avoiding cross-contamination from unrelated texts.
For experiments, we fine-tune various LLMs on standard instruction fine-tuning datasets and conduct bias-related experiments to evaluate the potential effect of packing on bias amplification. Additionally, we assess the impact of TFP on computational efficiency for SFT. Our findings indicate that TFP not only demonstrates superior performance across various models but also substantially lowers computational costs. The bias-related experiments show that TFP offers a flexible operational space, allowing for adjustments in the ratio of sensitive attributes (e.g., race) within packs to effectively manage bias.
@@ 2 Related Work
SFT and Alignment Fine-tuning is a prevalent strategy to enhance model performance on downstream tasks, evidenced in domains such as coding Wei et al. (2023); Luo et al. (2023) and arithmetic Xiang Yue (2023). Other work has highlighted the importance of consistency in format Liang et al. (2024), data quality Chung et al. (2022), and mixing tasks from different categories Longpre et al. (2023); Iyer et al. (2023) in SFT. As LLMs evolve, the risk of generating unsafe content increases Su et al. (2024); Wang et al. (2024a). Established methods for LLM alignment include instruction fine-tuning and reinforcement learning with human feedback (RLHF) Ouyang et al. (2022). Instruction fine-tuning, also known as SFT, refines pre-trained models using annotated instructional data, often preceding RLHF to aid initial alignment Touvron et al. (2023). RLHF employs reinforcement learning to adapt models based on feedback on generated responses. Although RLHF has been pivotal for developing systems like ChatGPT OpenAI (2021), isolated instruction fine-tuning can yield comparable outcomes Sun et al. (2023) with much less computational and labor costs.
Packing While packing is relatively less researched, it is a technique extensively used in frameworks like Hugging Face’s SFT Trainer  [^1] https://huggingface.co/docs/trl/sft_trainer to expedite inference and training. To prevent cross-contamination during self-attention calculation, existing packing approaches involve concatenating sequences into a single tensor and using masking to disregard elements from other sequences during computation Kundu et al. (2024). This method, including variations like LongAlign Bai et al. (2024) and Prepacking Zhao et al. (2024), enhances training efficiency and minimizes the cross-contamination impact on model performance. However, it necessitates calculating a distinct attention mask for each batch, complicating implementation and increasing memory consumption for masks, which can hinder the effectiveness of flash attention.
Our method differs from previous packing approaches that require masking to prevent cross-contamination. Instead, TFP creates packs of data points that provide useful context without the need for additional masking, simplifying implementation and reducing memory overhead.
@@ 3 Threshold Filtering Packing
The standard practice in packing is to form a pack by concatenating random samples until reaching the maximum context length Zhao et al. (2024). However, randomly concatenated packs do not provide additional learning signals and can lead to cross-contamination of sequences, compared to training on individual samples. In contrast, TFP generates more coherent packs by concatenating related and useful samples together, improving SFP performance and computational efficiency.
@@ 3.2 $k$-NN Packing
An intuitive method for packing is to apply $k$-NN and place each sample along with its retrieved top-$k$ neighbors in the same pack, referred to as $k$-NN packing. This approach maintains sample similarity within each pack but introduces a significant issue: data repetition. Some samples frequently appear as nearest neighbors for multiple others, leading to overlapping packs, i.e., $\exists i\neq j,P_{i}\cap P_{j}\neq\emptyset$. For instance, in the codealpaca2k dataset Chaudhary (2023), the sample "Construct a loop in Python to display all elements in a list" was included in 94 different packs with $k$-NN Packing, greatly reducing the diversity of pack content.
@@ 3.3 Threshold Filtering Packing Algorithm
To address these challenges, we propose packing data samples that provide meaningful context while avoiding repeated selection. A basic approach involves using a greedy algorithm to select samples with the smallest Euclidean distance to their corresponding embeddings, ensuring each sample is included only once. This is essentially a greedy algorithm for the TSP Applegate et al. (2006). Intuitively, $k$-NN can select the same data point multiple times when retrieving the nearest neighbors, whereas TSP selects each data point only once.
@@ 4 Experiment
Table 1: Comparison of different methods and Different Training Data: Alpaca, CodeAlpaca-2k, and GSM8K, with results represented by WR (Win Rate), HumanEval, and GSM8K, respectively. In the table, we follow the most common few-shot settings: using Win rate judged by PandaLM and a 0-shot setting for HumanEval, while the GSM8K task uses a 4-shot setting.
@@ 4.1 Experimental Setup
Datasets. We use commonly adopted datasets for instruction fine-tuning, which include tasks related to helpfulness, code-generation capabilities, and mathematical reasoning: (1) Alpaca dataset, generated from the Self-Instruct method Wang et al. (2023b) via the text-davinci-003 model Buruk (2023), covering various tasks such as arithmetic, coding, and question-answering. (2) Code Alpaca datasetChaudhary (2023), which aims to build and share an instruction-following LLaMA model for code generation.(3) GSM8K dataset Cobbe et al. (2021), curated to examine mathematical reasoning capabilities, comprises 8.8k high-quality arithmetic word problems designed at the grade school level.
For the fairness-related experiments, we use the Jigsaw Unintended Bias in Toxicity Classification task cjadams (2019) and Adult dataset Becker and Kohavi (1996). The Jigsaw Unintended Bias in Toxicity Classification task involves performing toxicity classification on comment texts published by the Civil Comments platform. It contains human-annotated demographic information such as race, gender, and religion. The goal is to ensure that models make predictions based on the text toxicity, rather than demographic information included in the text. We use race (Black and non-Black) as the protected attributes. The Adult dataset is a tabular dataset that includes 14 attributes of a person (e.g., age and education level) as input, to predict whether the person’s income exceeds \$50k per year. We evaluate the fairness of fine-tuned models based on the sensitive attribute of sex, specifically comparing “male” and “female”.
@@
Evaluation Metrics. We follow commonly used protocols Luo et al. (2024); Xiang Yue (2023); Ge et al. (2024) to evaluate SFT in LLMs. Specifically, we use PandaLM Wang et al. (2023a, 2024b) to evaluate the helpfulness of various models. PandaLM provides reproducible and automated comparisons between different LLMs. By providing PandaLM with the same context, it can compare the responses of different LLMs, offer reasons for the decisions, and provide a reference answer. We report the win rate (WR), which is the proportion of instances where the responses are favored over those produced by GPT-3.5 Brown et al. (2020). The code generation skills are enhanced using the Code Alpaca dataset Chaudhary (2023), while evaluation is conducted using the HumanEval dataset Chen et al. (2021). GSM8K dataset Cobbe et al. (2021) uses its own test set. We followed the most common few-shot settings: using Win rate judged by PandaLM and a 0-shot setting for HumanEval, while the GSM8K task uses a 4-shot setting.
We utilize the Llama2-7B Touvron et al. (2023), Llama3-8B AI@Meta (2024), and Mistral-7B Jiang et al. (2023) as the base LM in our experiments. Due to limited computation resources, we employ the QLoRA technique Dettmers et al. (2023) in all fine-tuning experiments. To ensure fair comparison, we maintain consistency in nearly all hyperparameters across all methods. For all results below, we run the experiments five times and report the mean and standard deviations for all compared methods.
@@ 4.2 Results
Table 2: Accuracy and group fairness metrics on Llama3-8B for the Jigsaw Dataset
@@
Table 3: Accuracy and group fairness metrics on Llama3-8B for the Adult Dataset
@@
Table 1 displays the results of fine-tuning on three downstream datasets: Alpaca Taori et al. (2023), codealpaca-2k Chaudhary (2023), and GSM8K Cobbe et al. (2021). We have the following key observations:
@@
Table 4: Training time (hrs) of LLaMA3-8B using different methods on GSM8K.
Our method, which packs similar data together in the same pack, may lead to imbalanced data across different demographic groups, potentially amplifying biases. We explore two approaches: TFP (balanced) and TFP (resampling) to achieve balance while maintaining text relevance. We also use balanced ratio and resampling methods on the data’s default sequence as controls. Particularly, inspired by DecodingTrust Wang et al. (2024a) that shows a balanced ratio across different groups in ICL can help improve LLM fairness, we investigate whether a balanced ratio of different demographic groups within a pack during SFT influences LLM fairness using classification tasks. We report prediction accuracy (ACC) and two common fairness metrics used for classification, equalized odds differences (eod) Hardt et al. (2016) and demographic parity difference (dpd) Zemel et al. (2013). Their detailed definitions can be found in Appendix A.
We use 0-shot and balanced 32-shot ICL settings (i.e., the ratio of different groups in ICL is 1:1) for evaluation following the experimental design of DecodingTrust Wang et al. (2024a), and we also explore the effects of a small number of samples with a balanced 4-shot setting. Experiments in Tables 2 and 3 were conducted on text and tabular datasets respectively, from which we have the following three observations.
(1) TFP (balanced) excels in fairness tasks for both text and tabular data, especially when the number of ICL examples is small. When the number becomes large (e.g., 32), the influence of the balanced ratio in ICL examples on fairness becomes dominant. The original TFP does not excel in fairness in 0-shot settings due to imbalanced data across social groups. TFP (resampling) results in instability due to the repeated sampling of certain data. By providing the model with hard negative examples (i.e. closely positioned samples with differing labels) within a pack, TFP enables more efficient learning from data with similar texts but different labels. Balancing the ratio within packs introduces many samples with similar texts but varying sensitive attributes, which helps mitigate biases in LLMs.
@@
Figure 2: Training time (hrs) on 2*L40S 48G and 1*3090 under different training methods.
4.3 Impact of Packing on Efficiency
@@
We observe that: (1) When data lengths are much shorter than the maximum token limit of LLMs, packing significantly reduces training costs, particularly in the fine-tuning phase, by optimizing CUDA matrix operations. This improvement is beneficial for both multi-GPU and single-GPU configurations. (2) Our packing method reduces the need for distinct attention masks per batch by allowing a single standard diagonal for the whole batch. It avoid increased mask memory consumption from incorrect implementation.

### v2

@@ Abstract
Packing for Supervised Fine-Tuning (SFT) in autoregressive models involves concatenating data points of varying lengths until reaching the designed maximum length to facilitate GPU processing. However, randomly concatenating data points can lead to cross-contamination of sequences due to the significant difference in their subject matter. The mainstream approaches in SFT ensure that each token in the attention calculation phase only focuses on tokens within its own short sequence, without providing additional learning signals for the preceding context. To address these challenges, we introduce Threshold Filtering Packing (TFP), a method that selects samples with related context while maintaining sufficient diversity within the same pack. Our experiments show that TFP offers a simple-to-implement and scalable approach that significantly enhances SFT performance, with observed improvements of up to 7% on GSM8K, 4% on HumanEval. Furthermore, results from bias benchmark datasets highlight TFP’s promising performance in improving fairness while also boosting prediction accuracy by 15%.
@@
Jiancheng Dong1, Lei Jiang1, Wei Jin2, Lu Cheng1 1University of Illinois Chicago, 2Emory University dongjiancheng77@gmail.com, {ljian43, lucheng}@uic.edu, wei.jin@emory.edu
@@ 1 Introduction
In Supervised Fine-Tuning (SFT) for large language models (LLMs), sequence lengths can vary substantially, requiring the wrapping of data into tensors to apply matrix operations optimized for CUDA and GPUs Raffel et al. (2020). As illustrated in Figure 1(a), vanilla fine-tuning pad shorter sequences with special tokens "[PAD]" up to the maximum sequence length. While this ensures uniformity, it introduces inefficiencies by including irrelevant padding tokens in the computation, wasting GPU resources and dilutes the model’s learning signal Kundu et al. (2024).
@@
However, randomly concatenating these data samples (Figure 1(b)), can result in sequence cross-contamination Krell et al. (2021). Cross-contamination occurs when predictions for one sequence are influenced by an unrelated sequence, complicating accurate predictions, especially when the subjects differ. For instance, imagine a model tasked with generating a multiplication table, followed by a prompt to list useful expressions in French. If these two sequences are concatenated without proper separation, the model might mix the tasks, producing results like "3 x 2 = Bonjour." This leads to incorrect and confusing outputs, where instructions and contexts become inappropriately blended. Moreover, current SFT pipelines Kundu et al. (2024) cause previous samples to provide no signal for predicting the next sample, thereby reducing learning efficiency and negatively impacting the few-shot performance of LLMs.
To address these challenges, in this work, we present Threshold Filtering Packing (TFP), a new packing approach that packs sequences of related yet diverse samples, encouraging context richness and reasoning across sample boundaries. Specifically, we employ a greedy algorithm inspired by the Traveling Salesman Problem (TSP) Applegate et al. (2006) to efficiently map out a path for segmentation into multiple packs. TFP further refine these packs by ensuring that overly similar samples are not grouped together. Setting the threshold is crucial in this context as it allows us to strike a balance between similarity and diversity within each pack, preventing homogeneity and ensuring the robustness of the generated packs. As shown in Figure 1(c), each sample is first converted to an embedding and then represented as a node in the graph. As shown in Figure 1(d), TFP employs threshold filtering to ensure each sample is distinct enough from recent ones, preventing the model from merely replicating previous outputs. This method results in packs that provide useful context while avoiding cross-contamination from unrelated texts.
For experiments, we fine-tune various LLMs on standard instruction fine-tuning datasets and conduct bias-related experiments to evaluate the potential effect of packing on fairness. Additionally, we assess the impact of TFP on computational efficiency for SFT. Our findings indicate that TFP demonstrates superior performance across various models. The bias-related experiments show that TFP offers a flexible operational space, allowing for adjustments in the ratio of sensitive attributes (e.g., race) within packs to effectively manage bias.
@@ 2 Related Work
SFT and Alignment Fine-tuning is a prevalent strategy to enhance model performance on downstream tasks, evidenced in domains such as coding Wei et al. (2023); Luo et al. (2024) and arithmetic Yue et al. (2024). Other work has highlighted the importance of consistency in format Liang et al. (2023), data quality Chung et al. (2022), and mixing tasks from different categories Longpre et al. (2023); Iyer et al. (2022) in SFT. As LLMs evolve, the risk of generating unsafe content increases Su et al. (2024); Wang et al. (2023a). Established methods for LLM alignment include instruction fine-tuning and reinforcement learning with human feedback (RLHF) Ouyang et al. (2022). Instruction fine-tuning, also known as SFT, refines pre-trained models using annotated instructional data, often preceding RLHF to aid initial alignment Touvron et al. (2023). RLHF employs reinforcement learning to adapt models based on feedback on generated responses. Although RLHF has been pivotal for developing systems like ChatGPT OpenAI (2021), isolated instruction fine-tuning can yield comparable outcomes Sun et al. (2023) with much less computational and labor costs.
Packing While packing is relatively less researched, it is a technique extensively used in frameworks like Hugging Face’s SFT Trainer  [^1] https://huggingface.co/docs/trl/sft_trainer to expedite inference and training. To prevent cross-contamination during self-attention calculation, existing packing approaches involve concatenating sequences into a single tensor and using masking to disregard elements from other sequences during computation Kundu et al. (2024). This method, including variations like LongAlign Bai et al. (2024) and Prepacking Zhao et al. (2024a), enhances training efficiency and minimizes the cross-contamination impact on model performance. However, it necessitates calculating a distinct attention mask for each batch, complicating implementation and increasing memory consumption for masks, which can hinder the effectiveness of flash attention.
In contrast, TFP avoids the need for masking to prevent cross-contamination. Instead, it forms data packs that provide relevant context without additional masking, streamlining implementation and reducing memory overhead.
@@ 3 Threshold Filtering Packing
The standard practice in packing is to form a pack by concatenating random samples until reaching the maximum context length Zhao et al. (2024a). However, randomly concatenated packs do not provide additional learning signals and can lead to cross-contamination of sequences, compared to training on individual samples. In contrast, TFP generates more coherent packs by concatenating related and useful samples together, improving SFT performance and computational efficiency.
@@ 3.2 $k$-NN Packing
An intuitive method for packing is to apply $k$-NN and place each sample along with its retrieved top-$k$ neighbors in the same pack, referred to as $k$-NN packing. This approach maintains sample similarity within each pack but introduces a significant issue: data repetition. Some samples frequently appear as nearest neighbors for multiple others, leading to overlapping packs, i.e., $\exists i\neq j,P_{i}\cap P_{j}\neq\emptyset$. For instance, in the CodeAlpaca dataset Chaudhary (2023), the sample "Construct a loop in Python to display all elements in a list" was included in 94 different packs with $k$-NN Packing, greatly reducing the diversity of pack content.
@@ 3.3 Threshold Filtering Packing Algorithm
To address these challenges, we propose packing data samples that provide meaningful context while avoiding repeated selection. Recent research on pretraining language models with related documents has inspired our approach Staniszewski et al. (2023); Shi et al. (2023); Zhao et al. (2024b), where similar documents are retrieved to enhance pretraining effectiveness. For Supervised Fine-Tuning datasets, we do not need to consider the degree of the starting node, as the shorter sequence length allows us to easily generate a complete graph. A basic approach involves using a greedy algorithm to select samples with the smallest Euclidean distance to their corresponding embeddings, ensuring that each sample is included only once. This is essentially a greedy algorithm for the TSP Applegate et al. (2006). Intuitively, $k$-NN can select the same data point multiple times when retrieving the nearest neighbors, whereas TSP ensures that each data point is selected only once.
@@ 4 Experiment
Table 1: Comparison of different methods and training datasets: Alpaca, CodeAlpaca, and GSM8K, with results represented by WR, HumanEval, and GSM8K, respectively. In the table, we follow the most common few-shot settings: using Win rate judged by PandaLM and a 0-shot setting for HumanEval, while the GSM8K task uses a 4-shot setting.
@@ 4.1 Experimental Setup
Datasets. We use commonly adopted datasets for instruction fine-tuning, which include tasks related to helpfulness, code-generation capabilities, and mathematical reasoning: (1) Alpaca dataset, generated from the Self-Instruct method Wang et al. (2023c) via the text-davinci-003 model Buruk (2023), covering various tasks such as arithmetic, coding, and question-answering. (2) CodeAlpaca datasetChaudhary (2023), which aims to build and share an instruction-following LLaMA model for code generation. (3) GSM8K dataset Cobbe et al. (2021), curated to examine mathematical reasoning capabilities, comprises 8.8k high-quality arithmetic word problems designed at the grade school level.
For the fairness-related experiments, we use the Jigsaw Unintended Bias in Toxicity Classification task Adams et al. (2019) and Adult dataset Becker and Kohavi (1996). The Jigsaw Unintended Bias in Toxicity Classification task involves performing toxicity classification on comment texts published by the Civil Comments platform. It contains human-annotated demographic information such as race, gender, and religion. The goal is to ensure that models make predictions based on the text toxicity, rather than demographic information included in the text. We use race (Black and non-Black) as the protected attributes. The Adult dataset is a tabular dataset that includes 14 attributes of a person (e.g., age and education level) as input, to predict whether the person’s income exceeds \$50k per year. We evaluate the fairness of fine-tuned models based on the sensitive attribute of sex, specifically comparing “male” and “female”.
@@
Evaluation Metrics. We follow commonly used protocols Luo et al. (2024); Yue et al. (2024); Ge et al. (2024) to evaluate SFT in LLMs. Specifically, we use PandaLM Wang et al. (2023b, 2024) to evaluate the helpfulness of various models. PandaLM provides reproducible and automated comparisons between different LLMs. By providing PandaLM with the same context, it can compare the responses of different LLMs, offer reasons for the decisions, and provide a reference answer. We report the win rate (WR), which is the proportion of instances where the responses are favored over those produced by GPT-3.5 Brown et al. (2020). The code generation skills are enhanced using the CodeAlpaca dataset Chaudhary (2023), while evaluation is conducted using the HumanEval dataset Chen et al. (2021). GSM8K dataset Cobbe et al. (2021) uses its own test set. We followed the most common few-shot settings: using Win rate judged by PandaLM and a 0-shot setting for HumanEval, while the GSM8K task uses a 4-shot setting.
We utilize the Llama2-7B Touvron et al. (2023), Llama3-8B AI@Meta (2024), and Mistral-7B Jiang et al. (2023) as the base LLM in our experiments. Due to limited computation resources, we employ the QLoRA technique Dettmers et al. (2023) in all fine-tuning experiments. To ensure fair comparison, we maintain consistency in nearly all hyperparameters across all methods. For all results below, we run the experiments five times and report the mean and standard deviations for all compared methods.
@@ 4.2 Results
Table 2: Accuracy and group fairness metrics on Llama3-8B for the Jigsaw Dataset with balanced few-shots.
@@
Table 3: Accuracy and group fairness metrics on Llama3-8B for the Adult Dataset with balanced few-shots.
@@
Table 1 displays the results of fine-tuning on three downstream datasets: Alpaca Taori et al. (2023), CodeAlpacaChaudhary (2023), and GSM8K Cobbe et al. (2021). We have the following key observations:
@@
Table 4: Training time (hrs) on 2*L40S and 1*3090 under different training methods.
Our method, which groups similar data together into packs, can sometimes lead to imbalanced representation across demographic groups, potentially amplifying biases. To address this, we explore two approaches: TFP (balanced) and TFP (resampling), both designed to achieve balance while maintaining text relevance. TFP (balanced) ensures that data with different sensitive attributes are evenly distributed within each pack, continuing this process until samples from one demographic group are exhausted. In contrast, TFP (resampling) addresses imbalances by resampling underrepresented data, ensuring all packs have a balanced representation of attributes, even when certain categories have fewer samples. For comparison, we apply balanced ratio and resampling methods to the data’s default sequence as controls, referred to as Balancing ratio and Resampling, respectively.
Inspired by DecodingTrust Wang et al. (2023a), which demonstrates that a balanced ratio across groups in ICL can improve LLM fairness, we examine whether maintaining a balanced ratio of demographic groups within packs during SFT affects LLM fairness in classification tasks. We report prediction accuracy (ACC) and two common fairness metrics used in classification: equalized odds difference (eod) Hardt et al. (2016) and demographic parity difference (dpd) Zemel et al. (2013). Detailed definitions of these metrics are provided in Appendix A. We use 0-shot and balanced 32-shot ICL settings for evaluation following the experimental design of DecodingTrust, and we also explore the effects of a small number of samples with a balanced 4-shot setting. Experiments in Tables 2 and 3 were conducted on text and tabular datasets respectively, from which we have the following three observations.
(1) TFP (balanced) excels in fairness tasks for both text and tabular data, especially when the number of balanced ICL few shot examples is large. We believe this is due the compounded effect of both balanced group ratio within the pack and the ICL demonstrations. The original TFP does not excel in fairness in 0-shot settings due to imbalanced data across social groups. TFP (resampling) results in instability due to the repeated sampling of same samples. By providing the model with hard negative examples (i.e. closely positioned samples with differing labels) within a pack, TFP enables more efficient learning from data with similar texts but different labels. Balancing the ratio within packs introduces many samples with similar texts but varying sensitive attributes, which helps mitigate biases in LLMs.
@@
Figure 2: Training time (hrs) of LLaMA3-8B using different methods on GSM8K.
Impact of Packing on Efficiency
@@
We observe that: (1) When data lengths are much shorter than the maximum token limit of LLMs, packing significantly reduces training costs, particularly in the fine-tuning phase, by optimizing CUDA matrix operations. This improvement is beneficial for both multi-GPU and single-GPU configurations. (2) Our packing method reduces the need for distinct attention masks per batch by allowing a single standard diagonal for the whole batch. It avoid increased mask memory consumption from incorrect implementation. (3) The time costs of graph traversal and pairwise similarity computation in our current experiments account for only 1% to 3% of the total fine-tuning process. Additionally, compared to training, these operations consume significantly less GPU memory.